OpenStudy (emmynimmy):

Find the equation of the line using the given information. The slope is undefined and it passes through the point (x, y) = (7, 4).

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Hint: A vertical line has no slope.

OpenStudy (skullpatrol):

Note: No slope is not the same as 0 slope.

OpenStudy (skullpatrol):

No slope means the slope is undefined. 0 slope means the slope has a value of 0.

OpenStudy (agent0smith):

Undefined means a vertical line, and is always of the form x= Here's an example: A vertical line passes through the point (5, 17) The equation of the line must be x=5. And another example: A vertical line passes through the point (-7, 6) The equation of the line must be x=-7.

OpenStudy (emmynimmy):

i put x=-4 and it says incorrect then it says Recall that the slope of a horizontal line is 0 and the slope of a vertical line is undefined. Note that the point-slope and slope-intercept formulas do not apply to vertical lines since the slope is undefined. If a line is vertical, which coordinate remains the same for every point on the line, the x-coordinate or the y-coordinate? Is the equation of a vertical line through the point (a, b); x = a, x = b, y = a, or y = b?

OpenStudy (agent0smith):

Look at my examples above, then find the equation of yours: A vertical line passes through the point (7, 4) equation of the line...?

OpenStudy (emmynimmy):

7?

OpenStudy (agent0smith):

That isn't an equation. See my examples and how i wrote the equation.

OpenStudy (emmynimmy):

x=7

OpenStudy (agent0smith):

Yes.

OpenStudy (emmynimmy):

okay, so i have this one/ Find the equation of the line using the given information. The slope equals zero and it passes through the point (x, y) = (1, −9).

OpenStudy (emmynimmy):

would it be that x=-9?

OpenStudy (agent0smith):

No, slope of zero is not the same as undefined. Slope of zero means the equation is y= Eg. The slope equals zero and it passes through the point (-9, 5) The equation must be y=5.

OpenStudy (emmynimmy):

so it would be y=1

OpenStudy (agent0smith):

Look at my example again.

OpenStudy (emmynimmy):

I'm confused.

OpenStudy (agent0smith):

Eg. The slope equals zero and it passes through the point (-15, 12) The equation must be y=12.

OpenStudy (emmynimmy):

y=-9

OpenStudy (agent0smith):

Yes.

OpenStudy (emmynimmy):

Find the slope of the line that passes through the given points. (−9, 2) and (5, −3) then it says Recall that the slope of a line is the rise over the run. Which coordinates tell the rise, the change in height? To find the run, should the x-coordinates be added or subtracted? Does it matter in which order the points are used?

OpenStudy (agent0smith):

For two points, \(\large (x_1, y_1)\) and \(\large (x_2, y_2)\) the slope is:\[\large m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] Eg. for the points \(\large (3, 5)\) and \(\large (2, 8)\) \[\Large m = \frac{ 8-5 }{ 2-3 } = \frac{ 3 }{ -1}=-3\]

OpenStudy (emmynimmy):

so m = -9

OpenStudy (emmynimmy):

no wait.

OpenStudy (agent0smith):

Show work.

OpenStudy (emmynimmy):

idk how you got 2,3 on the bottom

OpenStudy (emmynimmy):

I'm just confused.

OpenStudy (agent0smith):

Because those are the x2 and x1. On the top, you subtract the y values, on the bottom, you subtract the x values. (−9, 2) and (5, −3) \[\large m = \frac{ -3 - 2 }{ 5- (-9) }=?\]

OpenStudy (emmynimmy):

so on the top it would be -5 and on the bottom it would be -14?

OpenStudy (emmynimmy):

so it would be that divided or subtracted?

OpenStudy (agent0smith):

Two negatives make a positive. 5-(-9) = 5+9.

OpenStudy (emmynimmy):

14.

OpenStudy (emmynimmy):

so negative 5 and 14

OpenStudy (agent0smith):

\[\large m = \frac{ -3 - 2 }{ 5- (-9) }=\frac{ -5}{14 }\]You can't simplify any more.

OpenStudy (emmynimmy):

I'm trying to enter it into webasign and i can't see the side bar where it allows me to put a fraction there.

OpenStudy (agent0smith):

Then type that into a calculator

OpenStudy (emmynimmy):

-2.8

OpenStudy (emmynimmy):

nervmind got it.

OpenStudy (emmynimmy):

Find the slope of the line that passes through the given points. (−1, −2) and (3, 4)

OpenStudy (emmynimmy):

so slopes would be whats listed.

OpenStudy (agent0smith):

-5/14 is not -2.8

OpenStudy (emmynimmy):

i go it already

OpenStudy (agent0smith):

Just use the slope formula again

OpenStudy (emmynimmy):

3/2

OpenStudy (emmynimmy):

m=3/2

OpenStudy (agent0smith):

Yes

OpenStudy (emmynimmy):

A surveyor's map is drawn on a coordinate grid. Find the slope m of the line passing through the two labeled points. m =

OpenStudy (agent0smith):

Just use the slope formula again

OpenStudy (emmynimmy):

but i don't know what points for sure.

OpenStudy (emmynimmy):

wow wait I'm an idiot.

OpenStudy (emmynimmy):

okay

OpenStudy (agent0smith):

They're right there lol

OpenStudy (emmynimmy):

-6/11

OpenStudy (emmynimmy):

m=

OpenStudy (agent0smith):

Yes

OpenStudy (emmynimmy):

Find the x- and y-intercepts of the equation. −4x + 5y = 40 okay this ons confused me, i put none for x intercept and then 0,40 for y intercept but it says wrong

OpenStudy (agent0smith):

For the y-int: plug x=0 into the equation For the x-int: plug y=0 into the equation

OpenStudy (emmynimmy):

how do i do that?

OpenStudy (agent0smith):

plug x=0 means replace x with 0

OpenStudy (emmynimmy):

sides wouldn't be equal

OpenStudy (agent0smith):

Find the x- and y-intercepts of the equation. −4x + 5y = 40 For the y-int: plug x=0 into the equation: −4(0) + 5y = 40 solve for y

OpenStudy (emmynimmy):

y = 8

OpenStudy (agent0smith):

So (0, 8) is your y-int. Now find the x-int as i described above.

OpenStudy (emmynimmy):

-4, 0

OpenStudy (emmynimmy):

or 0,-4

OpenStudy (agent0smith):

No... Find the x- and y-intercepts of the equation. −4x + 5y = 40 For the x-int: plug y=0 into the equation: −4x + 5(0) = 40 solve for x

OpenStudy (emmynimmy):

-10

OpenStudy (agent0smith):

x-int is (-10, 0)

OpenStudy (emmynimmy):

Sketch a line with the given features. A y-intercept of (0, 5) and slope − 7 2

OpenStudy (emmynimmy):

slope = - 7/ 2

OpenStudy (emmynimmy):

its like this

OpenStudy (agent0smith):

I gotta go though :P Post it as a new question.